Find two functions f and g such that (f ∘ g)(x) = h(x).
h(x) = (3 − x)3
g(x) = ?
Hi, any help with this would be greatly appreciated.

Answer by ikleyn(52903)

.
Find two functions f and g such that (f ∘ g)(x) = h(x).
h(x) = (3 − x)^3                                                 <<<---===     I edited your post at this point.
g(x) = ?
~~~~~~~~~~~~~~~~~


            See how I edited your post to make sense from nonsense.


You can use  g(x) = (3-x);

             f(x) = x^3.


Then the composition  (f ∘ g)(x)  is  (3-x)^3,  exactly as requested.


    +-----------------------------+
    |   It is easy to understand. |
    +-----------------------------+


First function   g(x)  converts  x  into (3-x).


Second function  f(x)  converts  (3-x)  into  (3-x)^3.

Solved.